OpenStudy (anonymous):

What is the solution to the equation log 4 x – log 4 2 = 2 ? x = 16 x = 4 x = 32 x = 8

Parth (parthkohli):

\[\log_4 x - \log_4 2 \iff \log_4 \left(\dfrac{x}{2}\right)\]

Parth (parthkohli):

That leaves you with the following:\[\log_4\left(\dfrac{x}{2}\right) = 2\]Simply solve:\[\boxed{4^2 = \dfrac{x}{2}}\]

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

x=32

Parth (parthkohli):

Yes, 32.

OpenStudy (anonymous):

32 or 8????

OpenStudy (precal):